# Robustness of Neurosymbolic Reasoners on First-Order Logic Problems

### v1

@@ 1 Introduction
Figure 1: Illustration of our data and models. We test models in their ability to reason over default and counterfactual inputs, where key nouns were swapped (top). We compare fully neural models (LLMs) with neurosymbolic methods that combine LLMs with logical solvers. In our example the neural model fails on the counterfactual input but the neurosymbolic method makes correct predictions (bottom), suggesting higher robustness. Example taken from Wu et al. (2024).
LLMs have shown remarkable success on a range of different tasks including logical reasoning (Wei et al., 2022; DeepSeek-AI, 2024), mathematics (Lewkowycz et al., 2022), coding (Du et al., 2024), and creative tasks (Ramesh et al., 2021). These models have up to trillions of parameters and are trained to predict the most likely next word given the preceding words (Radford et al., 2018) on vast amounts of digitized data  (Brown et al., 2020; Chowdhery et al., 2023).
Wu et al. (2024) showed that these models suffer in test conditions that systematically differ from the inputs they observed during training by leveraging “counterfactual” (CF) tasks. CF tasks are carefully constructed to require the same reasoning as the original problem but with different assumptions. For example, for logical reasoning, Wu et al. (2024) demonstrated that replacing nouns and adjectives with less plausible alternatives in the input data to render the statements incompatible with any observed data can lead to a performance drop of 20%. This massive drop suggests that LLMs memorize their training data rather than learning how to reason logically. Figure˜1 illustrates this process.
@@
Our main research question is: are NS methods more robust to counterfactual variation than purely neural approaches? To answer this question, we employ LINC (Olausson et al., 2023), an NS method for logical reasoning that uses an LLM to translate NL into FOL, and compare against the LLM alone. Our experiments across LLMs with 7B to 32B parameters answer the question in the positive. To the best of our knowledge, we are the first to test the sensitivity of NS methods on counterfactual tasks.
However, we also find that LINC performs worse overall. We hypothesise that the model requires more explicit guidance to correctly convert NL statements to FOL. To address this, we then propose NSCoT which combines Chain-of-Thought (CoT) prompting Wei et al. (2022) and NS approaches to improve the overall performance of NS methods. Specifically, we include example CoT reasoning in the LLM’s in-context learning examples and prompt the LLM to generate its CoT reasoning when translating into FOL. We find that NSCoT substantially outperforms LINC but still lags behind a purely LLM approach with CoT.
@@ 2.1 Testing LLMs on perturbed data
Prior work has studied the sensitivity of LLMs to data perturbations. Jiang et al. (2024) demonstrated that simply replacing the primary noun in the prompt (e.g., from “Linda” to “Bob”) causes the model to fail, despite the logical structure of the task remaining unchanged. The paper concluded that LLMs suffer from the token bias problem, a phenomenon wherein LLMs exhibited a disproportionate reliance on frequently occurring lexical items such as specific nouns or structural cues to guide their reasoning process. Wu et al. (2024) introduced perturbations for a variety of tasks including arithmetic, code execution, logic, drawing, chord fingering, and chess. Their perturbations make the tasks deviate from the default, generally assumed conditions, which they called as “counterfactual” (CF). These CFs were manually constructed and carefully controlled to fix the difficulty levels of items and keep comparisons fair. They hypothesised that LLMs simply memorise their training examples rather than actually reasoning about problems. They found that although CoT reasoning and few-shot learning can reduce the gap in performance between default and CFs, a significant performance drop remains.
The token bias problem has also been studied in the mathematical domain. The GSM-Symbolic benchmark by Mirzadeh et al. (2025) systematically tests the impact of token bias by creating parsable templates and sampling different proper names and numerical values in mathematical problems. They showed that compared to default settings, numerical perturbations lead to about a 4% performance drop. A similar effect is found with proper names, further showing that the accuracy difference compounds when combined with the numerical perturbations.
Prior work has also shown that even when tasks remain within the reasoning capacity of humans, LLMs exhibit significant failures when problem complexity increases, presumably because such problems are rare in their pretraining data. For example, Shojaee et al. (2025) used the Tower of Hanoi problem as an example where they showed that while LLMs solve the problem with a small number of disks, their reasoning fails with larger number of disks.
Similar to prior work, we apply LLMs to a perturbed dataset to test their sensitivity to data perturbations. In contrast, however, we use LLMs in a neurosymbolic approach where we delegate the reasoning step to a symbolic solver. We note that in the literature, terms such as “counterfactual” are also used to describe hypothetical conditions that are false in the real world (Li et al., 2023). We use the term “counterfactual” hereinafter to be consistent with Wu et al. (2024), i.e., perturbed data samples.
@@ 2.2 Neurosymbolic reasoning
Recent work has explored the integration between LLMs and symbolic systems to improving the reasoning capabilities of LLMs. Such neurosymbolic methods introduce a two-stage pipeline where natural language is first translated into FOL statements, which are then passed into a symbolic solver for resolution. This positions the LLM to perform a more abstract role of semantic parsing rather than direct reasoning. For logical reasoning, recent neurosymbolic methods include LINC (Olausson et al., 2023), Logic-LM (Pan et al., 2023), and SatLM (Ye et al., 2023).
@@ 3.1 Dataset
We mainly work with the data from Wu et al. (2024), a subset of the FOLIO dataset Han et al. (2024) which has been turned into a ‘counterfactual’ data set. In FOLIO, ‘premises’ are paired with different ‘conclusions’ which either logically follow from the premises (True), or they don’t (False), or a conclusion cannot be drawn given the information in the premises (Uncertain). The task of a model is to classify the given the natural language premises and conclusion into one of these 3 labels.
Wu et al. (2024) manually swapped core noun variables in the premises with semantically implausible nouns which however do not alter the logical conclusion. Figure 1 shows an example. Intuitively, a robust reasoner would not be confused by this, while a brittle reasoning model which relies on surface cues would. There are 81 examples in this dataset, which we will refer to as RR (Reasoning or Reciting, from Wu et al.’s paper title).
@@ 3.2 Neurosymbolic Methods
As our neurosymbolic method, we use LINC (Olausson et al., 2023) where an LLM acts as a semantic parser to translate natural language premises and conclusions into FOL statements. These statements are then passed into a logic solver called Prover9 (McCune, 2005–2010) to predict the classification label. We use 8 in-context learning examples following Olausson et al. (2023).
The solver raises an error if an input cannot be parsed (i.e., if the LLM generate FOL statements that do not comply with Prover9’s format). To handle this, we follow Olausson et al. (2023): we prompt the LLM 10 times to obtain 10 generations, pass each of them to Prover9 to get a predicted label, and perform majority voting to get the final predicted label excluding the error cases. If all generations are errors, we count the prediction as wrong in performance evaluation.
@@
Table 1: Accuracies on the default and the counterfactual data (CF) as well as their differences ($\Delta$; 0 is best) on $R$. For a robust model, we expect a non-significant difference ($\Delta$) between the Default and CF condition. We mark brittle models for which this difference is significant ($p<0.05$; McNemar’s text McNemar (1947)) with an asterisk. The best result per model and metric is marked in bold.
@@ 3.3 Neural Approaches
We compare our neurosymbolic method LINC against three fully neural approaches of varying complexity following Olausson et al. (2023). The input prompt for each model contains 8 in-context learning examples followed by the given premises and conclusion to be evaluated:
@@
    Chain of Thought (CoT). Here, our 8-shot examples contain of premises, conclusion and the label, together with a human-created reasoning chain explaining why the label follows from the premises and conclusion pair. We use the reasoning chains given by Olausson et al. (2023). We lead the prompt with “Let’s think step by step”. The output consists of a generated reasoning chain and a final label.
@@ 3.4 Models
We test instruction-tuned open-source models from four families and of varying sizes: Mistral0.3 (Jiang et al., 2023), Qwen2.5 (Bai et al., 2023), Llama3.1 (Llama Team, 2024) and Gemma3 (Gemma Team, 2024). These span a representative set of models, with the Qwen and Mistral family are chosen for their focus on reasoning tasks while Llama, and Gemma signify more general-purpose language models. We test on model sizes between 7 billion and 32 billion parameters where available.
@@ 5.1 Method
We note here that the reasoning chain in each in-context learning example is generated using ChatGPT o3 model. This is different from the CoT approach we included in our baseline, where we included human-generated reasoning chains following Olausson et al. (2023). In addition, we prompt NSCoT to perform reasoning in response to each premise individually. Our CoT baseline performs reasoning over all premises at once, in one contiguous block.
@@ 5.2 Evaluation
We evaluate NSCoT on RR (N=81) under the same conditions as our main model (Table 1, right). In addition, we also validate NSCoT, LINC, and selected baseline methods on the full FOLIO validation data set of default premises (N=204). This is more than double the size of RR and contains examples that were excluded by Wu et al. (2024) in creating RR. Due to its increased size and diversity, we expect this dataset to be a more representative testbed for reasoning accuracy on default premises than RR. We report the accuracy numbers in Table˜2.
@@ 5.3 Results
Table 2: Accuracy on the FOLIO validation set (Han et al., 2024) which has more than double the size of $R$. M=Mistral, Q=Qwen, L=Llama, G=Gemma.
@@
From Table˜1, we observe that NSCoT outperforms (or is on par with) LINC on the default data of the smaller RR dataset only with Qwen2.5 32B and Llama 3.1. This is in contrast with the results we observe on the full FOLIO validation set of default premises in Table˜2 where NSCoT consistently outperforms LINC. It also conflicts with our results in Figure˜3 where NSCoT slightly outperforms LINC for all levels of complexity. We suspect that this discrepancy is due to the sample selection heuristics that Wu et al. (2024) used in creating RR and leave this investigation to future work. We contend that the results in Tables˜2 and 3 provide stronger evidence for the advantages of NSCoT due to the larger number and more complex examples in the full FOLIO validation set.
@@ 6.1 Class Distributions for Default vs Counterfactual Predictions
Figure˜5 compares the confusion matrices for Naive, CoT, NSCoT and LINC based on the Qwen7B instruction fine-tuned model and on the larger FOLIO validation data. We can observe that both the CoT and Naïve models show a higher False Negative (FN) rate for Uncertain class instances i.e. Naïve and CoT methods both tend to under-predict Uncertain. This suggests that neural methods overfit to surface regularities, confidently outputting categorical answers even when evidence is ambiguous. In contrast, the neurosymbolic methods (LINC and NSCoT) produce more Uncertain predictions. However, this comes with a trade-off: some of these are false positives because the LLM produces predicates with overlapping meaning (e.g., Dog and CuteDog), and Prover9 as a symbolic solver cannot detect this, thus predicting Uncertain for otherwise resolvable cases. Therefore, Uncertain sometimes reflects this kind of failures rather than truly uncertain cases.
@@ 6.2 Error analysis for NSCoT and LINC
To better understand the FOL conversion errors of both LINC and NSCoT, we manually classified the observed errors on all the examples from the FOLIO validation set where the methods generated FOL statements that did not comply with the Prover9 syntax. These were 341 cases for LINC and 366 cases for NSCoT, out of a total of 2040 queries (204 examples $\times$ 10 generations). We found two common classes of erros. The first are arity mismatches where predicates are used with inconsistent numbers of arguments across premises in the same instance (e.g. Likes(x,y) vs. Likes(x)). The second common error class pertains to unexpected tokens. This typically arises based on malformed or incomplete FOL strings (e.g., missing parentheses, unbalanced connectives), which cause Prover9 to throw parsing errors.
@@ 6.3 Faithfulness of CoT
The limits of CoT prompting have come under scrutiny by Paul et al. (2024) who showed that CoT rationales often reflect post-hoc justifications rather than the true decision process of the model. They modified the generated CoT reasoning such as making it incomplete, masking some tokens, or introducing some mistakes. They then re-queried the model with the modified reasoning and found that the model still gives the correct output, suggesting that the output does not actually depend on the printed reasoning chains, and that the reasoning chains are produced post-hoc after the final class prediction has already been computed.
@@ 7 Conclusion
Our work is limited to one type of reasoning namely first-order logic and should in the future be expanded to tasks such as math word problems (Huang et al., 2025) (using SymPy as a symbolic solver), coding, and planning. There is a need for systems which are more robust to variation and faithful under hard problems. Other limitations of this study include the small test data size and exclusion of larger-sized models due to computational resources.

### v2

@@ 1 Introduction
Figure 1: Illustration of our data and models. We test models in their ability to reason over default and counterfactual inputs, where key nouns were swapped (top). We compare fully neural models (LLMs) with neurosymbolic methods that combine LLMs with logical solvers. In our example the neural model fails on the counterfactual input but the neurosymbolic method makes correct predictions (bottom), suggesting higher robustness. Example taken from wu2024reasoningreciting.
LLMs have shown remarkable success on a range of different tasks including logical reasoning (wei2022cot; deepseek2024v3), mathematics (lewkowycz2022quantreasoning), coding (chen2021codeeval), and creative tasks (ramesh2021zeroshot). These models have up to trillions of parameters and are pretrained to predict the most likely next word given the preceding words (radford2018gpt) on vast amounts of digitized data  (brown2020fewshot; chowdhery2022palm). However, they cannot inherently perform formal rule based inference how a symbolic solver would.
wu2024reasoningreciting showed that these models suffer in test conditions that systematically differ from the inputs they observed during training by leveraging “counterfactual” (CF) tasks. CF tasks are carefully constructed to require the same reasoning as the original problem but with different assumptions. For example, for logical reasoning, wu2024reasoningreciting demonstrated that replacing nouns and adjectives with less plausible alternatives in the input data to render the statements incompatible with any observed data can lead to a performance drop of 20%. This massive drop suggests that LLMs memorize their training data rather than learning how to reason logically. Figure˜1 illustrates this process.
@@
Our main research question is: are NS methods more robust to counterfactual variation than purely neural approaches? To answer this question, we employ LINC (olausson2023linc), an NS method for logical reasoning that uses an LLM to translate NL into FOL, and compare against the LLM alone. Our experiments across LLMs with 7B to 32B parameters answer the question in the positive. To the best of our knowledge, we are the first to test the sensitivity of NS methods on counterfactual tasks.
However, we also find that LINC performs worse overall. We hypothesise that the model requires more explicit guidance to correctly convert NL statements to FOL. To address this, we then propose NSCoT which combines Chain-of-Thought (CoT) prompting wei2022cot and NS approaches to improve the overall performance of NS methods. Specifically, we include example CoT reasoning in the LLM’s in-context learning examples and prompt the LLM to generate its CoT reasoning when translating into FOL. We find that NSCoT substantially outperforms LINC but still lags behind a purely LLM approach with CoT.
@@ 2.1 Testing LLMs on perturbed data
Prior work has studied the sensitivity of LLMs to data perturbations. jiang2024tokenbias demonstrated that simply replacing the primary noun in the prompt (e.g., from “Linda” to “Bob”) causes the model to fail, despite the logical structure of the task remaining unchanged. The paper concluded that LLMs suffer from the token bias problem, a phenomenon wherein LLMs exhibited a disproportionate reliance on frequently occurring lexical items such as specific nouns or structural cues to guide their reasoning process. wu2024reasoningreciting introduced perturbations for a variety of tasks including arithmetic, code execution, logic, drawing, chord fingering, and chess. Their perturbations make the tasks deviate from the default, generally assumed conditions, which they called as “counterfactual” (CF). These CFs were manually constructed and carefully controlled to fix the difficulty levels of items and keep comparisons fair. They hypothesised that LLMs simply memorise their training examples rather than actually reasoning about problems. They found that although CoT reasoning and few-shot learning can reduce the gap in performance between default and CFs, a significant performance drop remains.
The token bias problem has also been studied in the mathematical domain. The GSM-Symbolic benchmark by mirzadeh2024gsmsymbolic systematically tests the impact of token bias by creating parsable templates and sampling different proper names and numerical values in mathematical problems. They showed that compared to default settings, numerical perturbations lead to about a 4% performance drop. A similar effect is found with proper names, further showing that the accuracy difference compounds when combined with the numerical perturbations.
Prior work has also shown that even when tasks remain within the reasoning capacity of humans, LLMs exhibit significant failures when problem complexity increases, presumably because such problems are rare in their pretraining data. For example, apple2025illusion used the Tower of Hanoi problem as an example where they showed that while LLMs solve the problem with a small number of disks, their reasoning fails with larger number of disks.
Similar to prior work, we apply LLMs to a perturbed dataset to test their sensitivity to data perturbations. In contrast, however, we use LLMs in a neurosymbolic approach where we delegate the reasoning step to a symbolic solver. We note that in the literature, terms such as “counterfactual” are also used to describe hypothetical conditions that are false in the real world (li2023counterfactual). We use the term “counterfactual” hereinafter to be consistent with wu2024reasoningreciting, i.e., perturbed data samples.
@@ 2.2 Neurosymbolic reasoning
Recent work has explored the integration between LLMs and symbolic systems to improving the reasoning capabilities of LLMs. Such neurosymbolic methods introduce a two-stage pipeline where natural language is first translated into FOL statements, which are then passed into a symbolic solver for resolution. This positions the LLM to perform a more abstract role of semantic parsing rather than direct reasoning. For logical reasoning, recent neurosymbolic methods include LINC (olausson2023linc), Logic-LM (pan2023logiclm), and SatLM (ye2023satlm).
@@ 3.1 Dataset
We mainly work with the data from wu2024reasoningreciting, a subset of the FOLIO dataset han-etal-2024-folio which has been turned into a ‘counterfactual’ data set. In FOLIO, ‘premises’ are paired with different ‘conclusions’ which either logically follow from the premises (True), or they don’t (False), or a conclusion cannot be drawn given the information in the premises (Uncertain). The task of a model is to classify the given the natural language premises and conclusion into one of these 3 labels.
wu2024reasoningreciting manually swapped core noun variables in the premises with semantically implausible nouns which however do not alter the logical conclusion. Figure 1 shows an example. Intuitively, a robust reasoner would not be confused by this, while a brittle reasoning model which relies on surface cues would. There are 81 examples in this dataset, which we will refer to as RR (Reasoning or Reciting, from wu2024reasoningreciting’s paper title).
@@ 3.2 Neurosymbolic Methods
As our neurosymbolic method, we use LINC (olausson2023linc) where an LLM acts as a semantic parser to translate natural language premises and conclusions into FOL statements. These statements are then passed into a logic solver called Prover9 (mccune2024prover9) to predict the classification label. We use 8 in-context learning examples following olausson2023linc.
The solver raises an error if an input cannot be parsed (i.e., if the LLM generate FOL statements that do not comply with Prover9’s format). To handle this, we follow olausson2023linc: we prompt the LLM 10 times to obtain 10 generations, pass each of them to Prover9 to get a predicted label, and perform majority voting to get the final predicted label excluding the error cases. If all generations are errors, we count the prediction as wrong in performance evaluation.
@@
Table 1: Accuracies on the default and the counterfactual data (CF) as well as their differences ($\Delta$; 0 is best) on $R$. For a robust model, we expect a non-significant difference ($\Delta$) between the Default and CF condition. We mark brittle models for which this difference is significant ($p<0.05$; McNemar’s test mcnemar1947note) with an asterisk. The best result per model and metric is marked in bold.
@@ 3.3 Neural Approaches
We compare our neurosymbolic method LINC against three fully neural approaches of varying complexity following olausson2023linc. The input prompt for each model contains 8 in-context learning examples followed by the given premises and conclusion to be evaluated:
@@
    Chain of Thought (CoT). Here, our 8-shot examples contain of premises, conclusion and the label, together with a human-created reasoning chain explaining why the label follows from the premises and conclusion pair. We use the reasoning chains given by olausson2023linc. We lead the prompt with “Let’s think step by step”. The output consists of a generated reasoning chain and a final label.
@@ 3.4 Models
We test instruction-tuned open-source models from four families and of varying sizes: Mistral0.3 (jiang2023mistral7b), Qwen2.5 (Bai_Bai_Chu_Cui_Dang_Deng_Fan_Ge_Han_Huang_etal._2023), Llama3.1 (dubey2024llama3) and Gemma3 (gemma2024). These span a representative set of models, with the Qwen and Mistral family are chosen for their focus on reasoning tasks while Llama, and Gemma signify more general-purpose language models. We test on model sizes between 7 billion and 32 billion parameters where available.
@@ 5.1 Method
We note here that the reasoning chain in each in-context learning example is generated using ChatGPT o3 model. This is different from the CoT approach we included in our baseline, where we included human-generated reasoning chains following olausson2023linc. In addition, we prompt NSCoT to perform reasoning in response to each premise individually. Our CoT baseline performs reasoning over all premises at once, in one contiguous block.
@@ 5.2 Evaluation
We evaluate NSCoT on RR (N=81) under the same conditions as our main model (Table 1, right). In addition, we also validate NSCoT, LINC, and selected baseline methods on the full FOLIO validation data set of default premises (N=204). This is more than double the size of RR and contains examples that were excluded by wu2024reasoningreciting in creating RR. Due to its increased size and diversity, we expect this dataset to be a more representative testbed for reasoning accuracy on default premises than RR. We report the accuracy numbers in Table˜2.
@@ 5.3 Results
Table 2: Accuracy on the FOLIO validation set (han-etal-2024-folio) which has more than double the size of $R$. M=Mistral, Q=Qwen, L=Llama, G=Gemma.
@@
From Table˜1, we observe that NSCoT outperforms (or is on par with) LINC on the default data of the smaller RR dataset only with Qwen2.5 32B and Llama 3.1. This is in contrast with the results we observe on the full FOLIO validation set of default premises in Table˜2 where NSCoT consistently outperforms LINC. It also conflicts with our results in Figure˜3 where NSCoT slightly outperforms LINC for all levels of complexity. We suspect that this discrepancy is due to the sample selection heuristics that wu2024reasoningreciting used in creating RR and leave this investigation to future work. We contend that the results in Tables˜2 and 3 provide stronger evidence for the advantages of NSCoT due to the larger number and more complex examples in the full FOLIO validation set.
@@ 6.1 Class Distributions for Default vs Counterfactual Predictions
Figure˜5 compares the confusion matrices for Naive, CoT, NSCoT and LINC based on the Qwen7B instruction fine-tuned model and on the larger FOLIO validation data. We can observe that both the CoT and Naïve models show a higher False Negative (FN) rate for Uncertain class instances i.e. Naïve and CoT methods both tend to under-predict Uncertain. This suggests that neural methods overfit to surface regularities, confidently outputting categorical answers even when evidence is ambiguous.
In contrast, the neurosymbolic methods (LINC and NSCoT) produce more Uncertain predictions. However, this comes with a trade-off: some of these are false positives because the LLM produces predicates with overlapping meaning, and Prover9 as a symbolic solver cannot detect this, thus predicting Uncertain for otherwise resolvable cases. Overlapping meaning refers to instances containing distinct predicates with shared denotation that form a disconnect in the logical flow. For example, if the LLM output contains the predicates “Dog” and “CuteDog” then Prover9 will not be able to be resolve them (i.e. that “CuteDog” implies “Dog”), causing the logical reasoning process to fail. A more elaborate example is included in Section˜A.2. The quantity of these errors is captured by the Uncertain false positives in the respective confusion matrices of LINC and NSCoT (Figure 5). This error class is an instance of the more general issue of implicit information in NL statements (e.g., that “CuteDog” implies “Dog”), which has also been noted in prior work olausson2023linc.
@@ 6.2 Error analysis for NSCoT and LINC
To better understand the FOL conversion errors of both LINC and NSCoT, we manually classified the observed errors on all the examples from the FOLIO validation set where the methods generated FOL statements that did not comply with the Prover9 syntax. Here, "error" is generations in which Prover9 was not able to resolve the given FOL statements due to the specified error classes. These were 341 cases for LINC and 366 cases for NSCoT, out of a total of 2040 queries (204 examples $\times$ 10 generations). We found two common classes of erros. The first are arity mismatches where predicates are used with inconsistent numbers of arguments across premises in the same instance (e.g. Likes(x,y) vs. Likes(x)). The second common error class pertains to unexpected tokens. This typically arises based on malformed or incomplete FOL strings (e.g., missing parentheses, unbalanced connectives), which cause Prover9 to throw parsing errors.
@@
We conducted preliminary tests with a verification module to refine generations which did not execute due to Prover9 errors on both methods. In this setup, the Prover9 error messages and a few examples of common syntax corrections were put into a new prompt, and the model was re-queried in a loop until the FOL expressions executed successfully or a maximum of 3 retries was reached. However we had similar findings to pan2023logiclm in which they showed that the execution rate of the symbolic prover increases using a refiner but at the same time the accuracy decreases due to more semantic errors.
@@ 6.3 Faithfulness of CoT
The limits of CoT prompting have come under scrutiny by lanham2023faithfulness who showed that CoT rationales often reflect post-hoc justifications rather than the true decision process of the model. They modified the generated CoT reasoning such as making it incomplete, masking some tokens, or introducing some mistakes. They then re-queried the model with the modified reasoning and found that the model still gives the correct output, suggesting that the output does not actually depend on the printed reasoning chains, and that the reasoning chains are produced post-hoc after the final class prediction has already been computed.
@@ 7 Conclusion
We note that further optimization is most critical for improving the accuracy of the neurosymbolic methods, rather than robustness, as the main source of performance degradation comes from inconsistent NL–FOL translations. Optimizing this stage would therefore allow the symbolic reasoning component compute on correct logical forms, yielding more faithful and accurate deductions.
@@
Our work is limited to one type of reasoning namely first-order logic and should in the future be expanded to tasks such as math word problems (huang2025mathperturb) (using SymPy as a symbolic solver), coding, and planning. There is a need for systems which are more robust to variation and faithful under hard problems. Other limitations of this study include the small test data size and exclusion of larger-sized models due to computational resources.